Noor Clothing House sells T-shirts. Each shirt costs TZS15 and is sold for TZS20 to 

customers. 

The maximum demand is 20,000 shirts per year and 

The average demand is 18,570 shirts per year

The average lead time is 57 days

the maximum lead time is 64 days.

Required: Compute the reorder level and the maximum, minimum level of stock of T-shirts for 


1
Expert's answer
2020-07-27T12:42:42-0400
"ATC= 15 TZS"

"p=20 TZS"


"Q_{max}=20000"

"Q_{av}=18570"

"t_{av}=57 days"

"t_{max}=64 days"

Reordering level= Maximum consumption × Maximum Reorder period



"Rl=20000\times 64=1280000"

Minimum stock level=Reordering level- (Normal level consumption × Normal Reorder period)


"Minimum=1280000-(18570\times57)=221510"

Maximum stock level= Reordering level+ Reordering quantity -(Minimum consumption × Minimum Reordering period)


"Q_{min}=Q_{av}-(Q_{max}-Q_{av})=18570-(20000-18570)=17140"


"t_{min}=t_{av}-(t_{max}-t_{av})=57-(64-57)=50"

"Max=1280000+18570-(17140 \times50)=441570"